A series of equal quarterly payments of $5000 for 12 years is equivalent to what present amount an interest rate of 9% compounded

a) Quarterly?

b) Monthly?

c) Continuously?

Please explain
Answer #1

Quarterly payment (PMT) = $ 5000

Number of years (nper) = 12 * 4 = 48 periods

a)

9% compounded quarterly

EAR = (1+9%/4)^4 -1 = 9.31 %

Present Value(PV) = =pv(9.31%/4,48,5000 PV(rate, nper, pmt, [fv], [type]) = $ 143,628. 20

b)

9% Compounded monthly

EAR = (1+ 9%/12)^12 -1 = 9.38%

Present Value = =PV(9.38%/4,48,-5000 PV(rate, nper, pmt, [fv], [type]) = $ 143,134.00

c)

9% Compounded continuously

EAR = e ^9% -1 = 9.41%

Present value = J =PV(9.41%/4,48,5000) PV(rate, nper, pmt, [fv], [type]) = $ 142,922.97